find the x and y components of a 400 N force at an angle of 125 to the x-axis
1
Expert's answer
2019-07-08T10:40:30-0400

This is the standard problem of the translation of polar coordinates "(r, \varphi) = (400, 125 \degree)" to rectangular "(x,y)". The relationship is

"x = r \cos \varphi = 400 \cos 125 \degree \approx -229.43, \\
y = r \sin \varphi = 400 \sin125 \degree \approx 327.66."